OpenStudy (anonymous):

What is the difference between a radical function and a radical equation? Briefly explain.

OpenStudy (anonymous):

an example for a radical function: \[f(x)=\sqrt(x)\] an example for a radical equation \[\sqrt(x+1)=3\]

OpenStudy (anonymous):

Thank you